Conferences in Research and Practice in Information Technology - Style Guide

John F. Roddick  Anne Other

School of Informatics and Engineering Flinders University of South Australia,

PO Box 2100, Adelaide, South Australia 5001,

Email: roddick@cs.flinders.edu.au

Abstract

This paper describes the manner in which papers should be formatted for papers adhering to the ACS series, Conference in Research and Practice in Information Technology. The abstract should be a maximum of 250 words and should clearly identify the content of the paper.

Keywords: As required..

1 Introduction

Normal text is 10 point, Times Roman in two columns. Page size is A4 with 0.8in borders on all sides, two columns with 0.2 in between them. There should not be a blank line between paragraphs. If the style is being used properly, a natural looking gap will be produced as L^{A}TEXthinks fit.

Headings should use the heading styles as shown. Numbering is automatic.

2 Heading Level 1

2.1 Heading Level 2

2.1.1 Heading Level 3

Headings below level 3 should be avoided. Tables and figures should ideally be confined to one column but where this is not possible should be located at the top of a page. Each be given a caption using the caption style, For example,

Figure Content

Figure 1: Caption

Citations should use the author date format only. For example,

...as proved by Snodgrass (1987) and Fayyed et al. (1996) and referred to in other works (Ben-Zvi 1982, Bentley 1986, Agrawal, Imielinski & Swami 1993) the process..., etc.

A basic, but much abused, rule for citations is that they should not be referred to as nouns. Eg. “(Snodgrass 1987) talks about” or “In (Snodgrass 1987), the process of ...” are both wrong. For further information regarding formats, please contact the editor-in-chief.

3 Submitting Camera Ready copy

Once your paper has been amended as required by the refereeing process, the source should be provided to the conference chair as required. Do not send a copy to CRPIT - this is the responsibility of the program chair alone. Note that only PDF and postscript files are the only acceptable formats. Do remember that CRPIT papers are A4, not US letter size (thus on the dvips you often have to specify -t a4) – see website.

Note that if errors in formatting are found they will need to be corrected before printing in extreme circumstances the paper will be left out of the volume. The most common errors in using this word template are:

1. Not having the correct margins they are 0.8 inch all round.

2. Not using A4 paper format when creating the pdf file.

4 References

References should be in the standard author-date format, examples of which are shown below (shown is a conference paper, a thesis, a book, a book section and a journal paper in that order). The required formats can be obtained by including the Harvard package and using style agsm. The files harvard.sty and agsm.bst are available at the CRPIT website.

References

Agrawal, R., Imielinski, T. & Swami, A. (1993), Mining association rules between sets of items in large databases, in ‘ACM SIGMOD International Conference on Management of Data’, Vol. 22, ACM Press, Washington DC, USA, pp. 207–216.

Ben-Zvi, J. (1982), The time relational model, Ph.D., University of California, Los Angeles.

Bentley, J. (1986), Programming pearls, Addison-Wesley.

You can use the \toappear to add a footnote for grant and other information. We prefer the copyright notice to be the last on the page.

Copyright \begin{matrix}c \\ ◯ \end{matrix}2006, Australian Computer Society, Inc. This paper appeared at Twenty-Ninth Australasian Computer Science Conference (ACSC2006), Hobart, Australia. Conferences in Research and Practice in Information Technology (CRPIT), Vol. 48. Vladimir Estivill-Castro and Gillian Dobbie, Ed. Reproduction for academic, not-for profit purposes permitted provided this text is included.

Fayyad, U. M., Piatetsky-Shapiro, G. & Smyth, P. (1996), From data mining to knowledge discovery: An overview, in U. Fayyad, G. Piatetsky-Shapiro, P. Smyth & R. Uthurusamy, eds, ‘Advances in Knowledge Discovery and Data Mining’, AAAI Press/MIT Press, pp. 1–34.

Snodgrass, R. (1987), ‘The temporal query language tquel’, ACM Transactions on Database Systems 12(2), 247–298.